I measured the transfer function of the AO path, and think that there are some features indicative of a problem somewhere in the IMC locking loop.
Regardless of the locking scheme used, high bandwidth control of the laser frequency relies on the fact that the laser frequency is slaved to the IMC cavity length with nearly zero error below ~50 kHz (assuming the IMC loop has a UGF > 100 kHz). In my single arm experiments, I didn't know what to make of the ripples that became apparent in the measured OLTF as the AO gain was ramped up.
Tonight, I measured the TF of the "AO path", which modifies the error point of the IMC, thereby changing the laser frequency.
Attachment #1 shows the result of the measurement.
I didn't use POX / POY as a sensor to confirm that this is real frequency noise, I will do so tomorrow. But it may be that realizing a stable crossover is difficult with so many features in the AO path.
Previous thread with a somewhat detailed characterization of the IMC loop electronics.